OpenStudy (anonymous):

I need help with a 6th grade algebra question involving patterns

OpenStudy (anonymous):

What is it?

OpenStudy (saifoo.khan):

Problems?

OpenStudy (anonymous):

What is it ?

OpenStudy (anonymous):

sigh

OpenStudy (anonymous):

Here's the question: The first 50 terms of a pattern sum to 2000. a) Describe the pattern. b) Describe how you found the sum. c) Give the 100th term of the pattern. d) List all the patterns that you found.

OpenStudy (anonymous):

anyone have some ideas?

OpenStudy (anonymous):

If it's addition you can formulate something like this: x + 2x + 3x + 4x...49x + 50x = 2000 The sum of numbers 1...n is obviously n * (n+1) / 2 in our case 50 * 51 / 2 = 1275. 1275x = 2000 x = ~1,5686274509803921568627450980392

OpenStudy (anonymous):

Or something simpler, like this: 50x = 2000 x = 40

OpenStudy (anonymous):

ok, so how do i turn x=40 into a pattern that the first 50 numbers = 2000?

OpenStudy (anonymous):

The latter meaning you just add 40 for 50 times: 40 + 40 + 40.... = 2000 So adding 40 100 times would be 4000.

OpenStudy (anonymous):

hmm, that works

OpenStudy (anonymous):

thank you